OpenStudy (anonymous):

int(int(sqrt(x^2+y^2), x = y .. sqrt(8*y)), y = 0 .. 8) Please help me! :( this should be calculated using polar cor

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

double inting on polars eh .... sounds delightful :)

OpenStudy (amistre64):

{SS} sqrt(x^2 +y^2) dr.dt

OpenStudy (amistre64):

be right back .....

OpenStudy (amistre64):

y <= x <= sqrt(8y) 0 <= y <= 8

OpenStudy (amistre64):

{SS} sqrt(x^2 +y^2) dx.dy seems more appropriate than dr.dt we gotta convert this to polar?

OpenStudy (anonymous):

yeah we need to convert it to polar,,what i dnt understand is..How I should change my boundaries

OpenStudy (amistre64):

id have to draw a picture..... :)

OpenStudy (amistre64):

i know the cap is a sphere ... but for the domain I need to see it

OpenStudy (anonymous):

This is the question though

OpenStudy (amistre64):

nice :)

OpenStudy (amistre64):

this is my picture

OpenStudy (amistre64):

as we sweep this across our dt doesnt really change does it?it goes for pi/4 to pi/2 if anything right?

OpenStudy (amistre64):

and our radius goes from 0 to the longest distance there is along the y=x line; or 8sqrt(2)

OpenStudy (anonymous):

yeah I dont understand what my radius should be..

OpenStudy (amistre64):

if i see it right; your radius goes from 0 to 8sqrt(2)

OpenStudy (anonymous):

how do u say 8sqart2..i dnt understand that part

OpenStudy (amistre64):

your radius is a constant change; your theta depends on sqrt(8y) right?

OpenStudy (amistre64):

ok... the radius is the distance from the origin to the end of the line right? it travels up the y=x line; which is a 45 degree angle the total length of that line is from 0, the origin, up to (8,8) the common point

OpenStudy (amistre64):

the slant of a rt tri with 45 degrees is sqrt(2); this goes 8 times longer... so: 0 <= radius <= 8sqrt(2)

OpenStudy (amistre64):

pi/4 <= theta <= sqrt(8y), just gotta figure out if that sqrt(8y) need to be altered...

OpenStudy (anonymous):

hello anthony sir,,,how are you ??????you are excellent.

OpenStudy (amistre64):

howdy sath :) im fine thnx

OpenStudy (anonymous):

i wish i were there with you!

OpenStudy (amistre64):

you just bored? or bad times?

OpenStudy (anonymous):

nah...but am great of yours.wanna learn something from you.

OpenStudy (anonymous):

atleast way to study.

OpenStudy (anonymous):

im trying to calculate it and I got 256pi/3...but it turns out to be wrong :(

OpenStudy (amistre64):

{SS} f(x,y) dA <=> {SS} f(rcos, rsin)r dr.dt

OpenStudy (anonymous):

so i have to integrate this right int(int(r^2, r = 0 .. 8*sqrt(2)), theta = (1/4)*Pi .. (1/2)*Pi)

OpenStudy (amistre64):

i think the upper limit for the theta is spose to correspond to the x = sqrt(8y) function somehow

OpenStudy (amistre64):

x = r cos(t) y = r sin(r) right r cos(t) = sqrt(8y) r cos(t) = sqrt(8 rsin(t))

OpenStudy (amistre64):

r^2 cos^2(t) = 8r sin(t) solve for 't' i think

OpenStudy (amistre64):

or solve for r :)

OpenStudy (amistre64):

finding 't' sounds more reasonable to me; unless we can convert around to make it simpler

OpenStudy (anonymous):

this is sooo difficult.. :(

OpenStudy (amistre64):

well the good news is that the equation sqrt(x^2 +y^2) is just the unit sphere :) and should be easy to convert to polars

OpenStudy (anonymous):

yeah..as I said earlier after I used those boundaries (theetha bound as pi/4to pi/2 and r bound 0 to \[8\sqrt{2}\] the answer I got was 256pi/3 but its wrong

OpenStudy (amistre64):

r^2 cos^2(t) = 8r sin(t) ; /r r cos^2(t) = 8 sin(t) ; indentity r (1-sin^2(t)) = 8 sin(t) r - r sin^2(t) = 8 sin(t)...... it gonna be wrong because you sweep 90 degrees each time thru. You have to only sweep out the are from p/4 to the curve x=sqrt(8y)

OpenStudy (amistre64):

you are ending up taking the area 1/16th of a sphere if you do pi/4 to pi/2. and we dont want the whole area...

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

so is the answer pi/16?

OpenStudy (amistre64):

i got no clue what the answer is yet :) I could check it with the x and y stuff, but the polar version I aint got ...

OpenStudy (anonymous):

yeah please...

OpenStudy (amistre64):

hold on...

OpenStudy (anonymous):

sure

OpenStudy (amistre64):

well, i had to seek a higher source lol

OpenStudy (anonymous):

but still it turns out to be incorrect..

OpenStudy (amistre64):

is it looking for an exact? or an approximation?

OpenStudy (anonymous):

I think we should keep the answers in terms of pi..

OpenStudy (amistre64):

i aint having no luck at it....

OpenStudy (anonymous):

its fine then..thanks a lot :)

OpenStudy (anonymous):

i have the answer exactly if you want. 1024/45+(1024/45)*sqrt(2)

OpenStudy (anonymous):

and i have the correct integral. I used maple to solve though. I can try to do it by hand if you want.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}\int\limits_{0}^{\frac{8\sin(\theta)}{\cos^{2}(\theta)}}r^{2}drd \theta\]

OpenStudy (anonymous):

can you read that?

OpenStudy (anonymous):

ya i can read it..

OpenStudy (anonymous):

omg its crct....pls tell me hw u do it

OpenStudy (anonymous):

i have it down to this: \[8^3/6*\int\limits_{0}^{\pi/4}x*\sqrt{1+x}dx\]

OpenStudy (anonymous):

ok did you draw the region?

OpenStudy (anonymous):

wait nevermind i made a mistake that simplified integral is wrong

OpenStudy (anonymous):

but the first one is right

OpenStudy (anonymous):

the picture is a little strange the curve should be y=(x^2)/8, but it doesnt make much difference

OpenStudy (anonymous):

so we know at the top limit: x=sqrt(8y) x^2=8y r^2cos^2(theta)=8*r*sin(theta) r=8sin(theta)/(cos^2(theta)) thats where the upper limit came from

OpenStudy (anonymous):

and as a lower limit r=0

OpenStudy (anonymous):

then the maximum r is 8sqrt(2) so we want to find corresponding theta 8sqrt(2)=8sin(theta)/cos^2(theta) sqrt(2)=sin(theta)/cos^2(theta) at Pi/4 sin(theta)=sqrt(2)/2 and cos^2(theta)=1/2 so Pi/4 is our solution. theta from 0 to Pi/4

OpenStudy (anonymous):

so you have that integral i sent you, and you need to solve that

OpenStudy (anonymous):

the first integration is standard, but the second is a bit messy.

OpenStudy (anonymous):

here is the correct final integral: \[8^{3}/2\int\limits_{0}^{1}x \sqrt{1+x}dx\]

OpenStudy (anonymous):

sorry should be 8^3/6

OpenStudy (anonymous):

you can do the final integral by parts

OpenStudy (anonymous):

if you want to know how I got to the integral above let me know

OpenStudy (anonymous):

oh can u please tell me how u say that the maximum r is 8sqrt2

OpenStudy (anonymous):

sure r is in the shaded region and the maximum is at the intersection of the two curves. there is a 45 45 90 right triangle with legs of size 8 so the hypotenuse is 8sqrt(2)

OpenStudy (anonymous):

sqrt(8^2+8^2)=sqrt(128)=sqrt(64*2)=8sqrt(2)

OpenStudy (anonymous):

does the rest of the initial set up make sense?

OpenStudy (anonymous):

Need few minutes more..Im looking at your solution

OpenStudy (anonymous):

when you do the first integration and plug in limits you get: \[\int\limits_{0}^{\pi/4}8^{3}/3*(\sin^{3}(\theta)/\cos^{6}(\theta)d \theta\]

OpenStudy (anonymous):

or: \[8^{3}/3\int\limits_{0}^{\pi/4}\tan^{3}(\theta)\sec^{3}(\theta)d \theta\]

OpenStudy (anonymous):

use substitution and chose tan^2(theta)=x 2tan(theta)sec^2(theta)=dx 1+tan^2(theta)=1+x sec^2(theta)=1+x sec(theta)=sqrt(1+x) putting that all together and changing the limits of integration: \[8^{3}/6\int\limits_{0}^{1}x \sqrt{1+x} dx\]

OpenStudy (anonymous):

now by parts choose u=x dv=(1+x)^1/2 du=1 v=2((1+x)^(3/2))//3 and then you have a final simple power rule integral to solve

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}\int\limits_{0}^{8\sqrt{2}}r^2 drd \Theta\]

OpenStudy (anonymous):

so what I have to do is solve this right

OpenStudy (anonymous):

sorry i didnt mean the upper limit of integration on r was 8sqrt(2), I meant that was the r that let you find the limits of theta. the upper limit on r=8sin(theta)/cos^2(theta)

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{8\sin(\theta)}{\cos^{2}(\theta)}}r^{2}dr\] thats the inner integral

OpenStudy (anonymous):

whole thing you need to solve: \[\int\limits\limits_{0}^{\pi/4}\int\limits\limits_{0}^{\frac{8\sin(\theta)}{\cos^{2}(\theta)}}r^{2}drd \theta\]

OpenStudy (anonymous):

theta from 0 to Pi/4 you can see if you draw the picture correctly

OpenStudy (anonymous):

oh i see.but like isnt solving the inner integral difficult?

OpenStudy (anonymous):

no its not bad. Did you see my posts above, I wrote out the steps

OpenStudy (anonymous):

let me check..I missed it i guess :) thanks a tons btw

OpenStudy (anonymous):

you're welcome :) i want to be a math professor so this is fun for me.

OpenStudy (anonymous):

btw hw dd u get 8^3/3 for r

OpenStudy (anonymous):

thats from integrating r^2 int(r^2)=r^3/3. then plug in the upper and lower limits for r the constant term is 8^3/3

OpenStudy (anonymous):

YEAH BUT THE upper and lower limits for r is 0 to sin(theeta)/cos^2(theta) right? that meas we have to substitue those two right?